OpenStudy (anonymous):

Solve the intial value problem. d^2r/dt^2=3/t^4; dr/dt t=1 =1, r(2)=2 Please Help!

OpenStudy (anonymous):

\[\frac{d^2r}{dt^2}=\frac{3}{t^4} \\ \frac{d}{dt}(\frac{dr}{dt})=\frac{3}{t^4} \\ hence \ \ by \ \ integration \\ \frac{dr}{dt}=-\frac{1}{t^3}+c\]

OpenStudy (anonymous):

am i clear?

OpenStudy (anonymous):

yes correct

OpenStudy (anonymous):

now u have dr/dt t=1 =1 put this bondary condition in the equation what will u get?

OpenStudy (anonymous):

i put the boundary into dr/dt correct?

OpenStudy (anonymous):

C=1?

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

now u have \[\frac{dr}{dt}=-\frac{1}{t^3}+1\] integrate again to get the final answer

OpenStudy (anonymous):

this is where i have a problem where i have to intergrate again

OpenStudy (anonymous):

1/3t^2+1t+C?

OpenStudy (anonymous):

1/2t^2+1t+C

OpenStudy (anonymous):

oh ok i eas really wrong

OpenStudy (anonymous):

i meant was not eas

OpenStudy (anonymous):

u have to integrate until u have such terms that include derivatives of your unknown function

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

what is final answer?

OpenStudy (anonymous):

so is the answer 1/2t^2+1t-1/8 ?

OpenStudy (anonymous):

yes thats completely right

OpenStudy (anonymous):

ok thanks so much for your help =)

OpenStudy (anonymous):

welcome